OpenStudy (anonymous):

Solve the system of equations x + 3y = 32 -3x + 2y = 3

OpenStudy (luffingsails):

Hi bart, A good approach to these is to solve one of the equations for one of the variables. So, if I solved the first equation for x, I would get: x = 32 - 3y I can then put this into the second equation and solve. -3(32 - 3y) + 2y = 3 -96 + 9y + 2y = 3, after distribution 11y = 99, after some simplification y = 9, after some more. Hope that helps.

OpenStudy (anonymous):

\[x + 3y = 32\]\[-3x + 2y = 3\] \[x=32-3y\]\[-3(32-3y)+2y=3\] \[x=32-3y\]\[-96+9y+2y=3\] \[x=32-3y\]\[11y=99\] \[x=32-3y\]\[y=9\] \[x=32-3\times9\]\[y=9\] \[x=5\]\[y=9\]